# Introducing FOReCAst: The Future Outcome Reasoning and Confidence Assessment Benchmark

### v3

@@ 5.2 Results, Findings, and Discussions
Table 4 shows that while some models reach reasonable accuracy on binary classification tasks, their confidence calibration remains inconsistent. For instance, GPT2-XL achieves a relatively strong accuracy of 0.67 but has a higher Brier score of 0.45, whereas OLMo-7B-Instruct achieves a similar accuracy (0.65) with a lower Brier score of 0.38, indicating better alignment between predicted probabilities and actual outcomes. This discrepancy suggests that models may differ not only in their ability to identify the correct answer, but also in how they distribute confidence over binary options.
Importantly, model performance does not appear to correlate straightforwardly with scale or training recency. For example, LLaMA-7B, a more recent model, shows lower accuracy (0.57) and higher Brier score (0.52) than older models like GPT2-XL. This supports the observation that accuracy in binary forecasting tasks depends on a combination of world knowledge, probabilistic reasoning, and how well models can handle epistemic uncertainty, particularly when questions are unresolved or partially underspecified at prompt time. The gap between classification accuracy and calibration further underscores the importance of evaluating both, as good classification alone may mask unreliable probability estimates.
Quantity Estimation Timeframe Prediction
Model N APE ($\downarrow$) MAE ($\downarrow$) CRPS ($\downarrow$) Model N ADE ($\downarrow$) CRPS ($\downarrow$)
GPT2 81 0.23 0.87 0.86 GPT2 26 0.99 0.99
GPT2-XL 81 0.03 0.75 0.72 GPT2-XL 26 1.00 1.00
Pythia-14m 77 0.17 0.88 0.85 Pythia-14m 25 1.00 1.00
Pythia-160m 77 0.03 0.78 0.76 Pythia-160m 25 1.00 1.00
Pythia-2.8b 77 0.07 0.81 0.79 Pythia-2.8b 25 0.97 0.96
Bloom-560m 43 0.05 0.73 0.71 Bloom-560m 12 1.00 1.00
Bloom-7b1 43 0.06 0.75 0.72 Bloom-7b1 12 1.00 1.00
Llama-7b 32 0.06 0.64 0.61 Llama-7b 10 0.98 0.98
OLMo-1B 23 0.21 0.79 0.76 OLMo-1B 6 1.00 1.00
OLMo-7B 23 0.22 0.84 0.82 OLMo-7B 6 1.00 1.00
OLMo-7B-Instruct 23 0.24 0.80 0.77 OLMo-7B-Instruct 6 1.00 1.00
OLMo-2-7B 20 0.02 0.65 0.62 OLMo-2-7B 4 1.00 1.00
OLMo-2-7B-Instruct 20 0.09 0.60 0.57 OLMo-2-7B-Instruct 4 0.87 0.82
Table 5: Performance of forecasting models on Quantity Estimation (left) and Timeframe Prediction (right) tasks in FOReCAst. All metrics are lower-is-better. APE: Absolute Percentage Error, MAE: Mean Absolute Error, CRPS: Continuous Ranked Probability Score, ADE: Absolute Days Error.
@@
Table 5 demonstrates that timeframe prediction—estimating when an event will occur—is especially difficult for current models. Across the board, both ADE and CRPS values remain close to 1.0, indicating little improvement over naive or uninformed baselines. Only OLMo-2-7B-Instruct achieves a CRPS noticeably below this range (0.82), but even this is far from well-calibrated. The generally poor performance suggests that temporal forecasting presents challenges beyond simple factual retrieval or classification.
This difficulty likely reflects the inherent complexity of temporal prediction, which often requires models to reason over event dynamics, causality, and elapsed time—all aspects that are underrepresented in standard pretraining corpora. In contrast to Boolean tasks, where resolution criteria are clearly defined and outcomes are binary, timeframe questions require mapping language input to a continuous temporal target, which demands a finer-grained understanding of how events evolve over time. That most models fail to depart meaningfully from baseline-level CRPS indicates that LLMs may lack sufficient temporal inductive biases or exposure to structured time-referenced reasoning.
@@
As shown in Table 5, quantity estimation—predicting a numerical outcome—exhibits a broader range of performance across models. Some models such as Pythia-160M and GPT2-XL achieve very low absolute percentage error (APE = 0.02 and 0.03 respectively), indicating strong point prediction capability. However, their CRPS values (0.76 and 0.72) suggest relatively poor calibration. In contrast, OLMo-2-7B-Instruct, with a higher APE of 0.08, attains the lowest CRPS (0.57), suggesting it better captures uncertainty even if its point estimates are less precise.
This divergence illustrates that point prediction accuracy and confidence calibration are not necessarily aligned. Predicting a correct numerical value does not guarantee that a model’s probability distribution around its prediction reflects uncertainty appropriately. Since CRPS penalizes both distance from the true value and over/under-confidence, this metric provides a more comprehensive view of performance in numerical forecasting tasks. These results suggest that even relatively strong predictive models may require additional mechanisms—such as structured uncertainty modeling or calibration-sensitive objectives—to produce well-calibrated forecasts in quantity estimation settings.
@@
Model size alone does not reliably predict forecasting performance across tasks. Within the Pythia family, Pythia-2.8b occasionally improves over smaller variants in certain metrics—for example, it achieves lower CRPS than Pythia-14m on quantity estimation—but these gains are neither large nor consistent. In some cases, smaller models like Pythia-160M outperform their larger counterparts in point prediction metrics such as APE. A similar trend appears within the OLMo family: while the OLMo-2-7B variants perform competitively in calibration metrics (e.g., CRPS), this is not accompanied by uniform improvements in accuracy or absolute error.
These observations suggest that increases in parameter count do not automatically translate to improved forecasting performance. While larger models often benefit from greater representational capacity, the forecasting task—particularly when it involves expressing well-calibrated uncertainty—may place demands that are orthogonal to model size. Specifically, forecasting tasks may require targeted reasoning over time, quantities, or epistemic uncertainty, which are not clearly emphasized during standard pretraining or scaled parameter growth. These findings indicate that effective forecasting may depend more on model behavior during decoding or task-specific adaptation, rather than scale alone.
Model Accuracy ($\uparrow$) F1 ($\uparrow$) Brier ($\downarrow$) ADE ($\downarrow$) CRPS (T) ($\downarrow$) APE ($\downarrow$) MAE ($\downarrow$) CRPS (Q) ($\downarrow$)
@@
Table 6 compares base and instruct-tuned variants of OLMo-7B and OLMo-2-7B across Boolean, timeframe, and quantity forecasting tasks. For Boolean questions, OLMo-7B-Instruct achieves higher accuracy (0.65 vs. 0.21) and a lower Brier score (0.38 vs. 0.82), indicating better confidence calibration. In timeframe prediction, the instruct-tuned OLMo-2-7B-Instruct improves uncertainty estimation, with an ADE of 0.87 and CRPS of 0.82, compared to 0.9981 and 0.9970 for the base OLMo-2-7B. For quantity estimation, instruct-tuned models have slightly higher APE but lower MAE and CRPS, suggesting better uncertainty calibration. While these results point to potential benefits of instruction tuning for probabilistic reasoning, the sample size remains limited due to the recency of instruction-tuned models, and the observed gains may not be statistically robust.
Aggregation Method N Accuracy ($\uparrow$) F1 ($\uparrow$) Brier ($\downarrow$)
@@
Table 7: Ablation study on different aggregation methods for extracting predictions and confidence from Llama-7B outputs.
@@
Table 7 compares different aggregation methods for deriving the final prediction and confidence estimate from the top 10 outputs of Llama-7B. Bayesian Aggregation achieves the highest accuracy (0.58) and F1 score (0.35), suggesting it is the most effective at identifying correct outcomes. However, Weighted Average yields a significantly lower Brier score (0.29), indicating superior confidence calibration compared to others.
Majority Vote, Highest Confidence, and Logit Mean Probability produce comparable accuracy and F1 scores but have noticeably higher Brier scores, suggesting weaker uncertainty estimation. These results highlight that even when point prediction performance is similar, aggregation methods substantially impact confidence reliability. The challenge remains in developing techniques that optimize both accuracy and calibration simultaneously, emphasizing the importance of uncertainty-aware forecasting. In preliminary experiments, we also explored verbalized confidence, prompting the model to explicitly generate a probability (e.g., “I’m 80% confident…”). However, most models either failed to produce enough meaningful numeric estimates or generated inconsistent responses, particularly on harder questions. This highlights a limitation in current instruction-following capabilities when it comes to quantifying uncertainty directly.
A promising direction for future work is to fine-tune models specifically to produce structured confidence outputs, possibly by supervising against calibrated targets or training with ranking-based objectives. Additionally, integrating task-specific priors (e.g., event resolution frequency, temporal proximity) into aggregation could further improve calibration. Hence, joint modeling of correctness and uncertainty remains a key open challenge for real-world deployment of forecasting systems.
@@
This observation suggests that forecasting is meaningfully different from retrieval-based tasks such as fact-checking or question answering. While those tasks often benefit from training data that includes the target facts or related content, forecasting involves reasoning under uncertainty about events that had not yet occurred during training. Models are not explicitly trained to perform such reasoning, and pretraining objectives generally do not reward predictions about future developments. The absence of a correlation between cutoff proximity and performance highlights that forecasting requires capabilities beyond knowledge access, including abstraction, probabilistic reasoning, and temporal inference. These factors make the task more complex and reinforce the need for dedicated methods that go beyond static knowledge retrieval.
@@ 7 Limitations
FOReCAst is constructed entirely from Metaculus, a platform known for its high-quality, well-resolved forecasting questions and active user base. This makes it a strong and consistent foundation for developing the task of probabilistic forecasting with LLMs. However, it does not cover the full diversity of forecasting practices, such as different question formats, domain emphases, or forecasting styles seen in other platforms like Good Judgment Open or domain-specific expert settings. In this work, our primary focus is to define the task rigorously and establish a benchmark that is methodologically clean and scalable. Expanding to additional data sources and platforms is part of our future work and should help test generalization across forecasting cultures and environments.
Our method for constructing gold confidence labels is based on community-aggregated forecasts, which offer a practical and scalable proxy for crowd belief and uncertainty. While this approach enables systematic evaluation across hundreds of questions, it may not fully capture the variability and richness of individual forecasters’ reasoning or levels of disagreement. Moreover, the aggregated predictions may sometimes be influenced by group dynamics or correlated sources of error. We acknowledge these trade-offs, and view our current approach as a reasonable starting point given the limited availability of high-quality, large-scale confidence annotations. Future work could explore incorporating richer uncertainty representations, such as forecast distributions, forecaster rationales, or disagreement metrics.
Finally, our dataset is currently limited to English-language questions and forecasts, reflecting the language used by Metaculus and many existing NLP resources. While this is a reasonable initial scope for establishing the task, we recognize that forecasting behavior, uncertainty expression, and interpretation of probabilities may vary significantly across linguistic and cultural contexts. Extending FOReCAst to include multilingual questions and community forecasts is a promising direction for improving the inclusiveness and robustness of future benchmarks. We also note that some dimensions, such as temporal adaptation and real-time reasoning, remain out of scope here but are essential long-term goals for building forecasting-capable systems.
@@ Appendix E Additional Results
Model Accuracy ($\uparrow$) F1 ($\uparrow$) Brier ($\downarrow$) ADE ($\downarrow$) CRPS (T) ($\downarrow$) APE ($\downarrow$) MAE ($\downarrow$) CRPS (Q) ($\downarrow$)
@@
OLMo-2-7B-Instruct 0.59 0.42 0.40 0.84 0.83 0.13 0.69 0.67
Table 10: Combined forecasting performance for cutoff 2017-12-01. CRPS (T) denotes the Continuous Ranked Probability Score for Timeframe Prediction, while CRPS (Q) denotes the Continuous Ranked Probability Score for Quantity Estimation.
Model Accuracy ($\uparrow$) F1 ($\uparrow$) Brier ($\downarrow$) ADE ($\downarrow$) CRPS (T) ($\downarrow$) APE ($\downarrow$) MAE ($\downarrow$) CRPS (Q) ($\downarrow$)
@@
OLMo-2-7B-Instruct 0.59 0.41 0.40 0.83 0.82 0.12 0.70 0.68
Table 11: Combined forecasting performance for cutoff 2020-03-01. CRPS (T) denotes the Continuous Ranked Probability Score for Timeframe Prediction, while CRPS (Q) denotes the Continuous Ranked Probability Score for Quantity Estimation.
Model Accuracy ($\uparrow$) F1 ($\uparrow$) Brier ($\downarrow$) ADE ($\downarrow$) CRPS (T) ($\downarrow$) APE ($\downarrow$) MAE ($\downarrow$) CRPS (Q) ($\downarrow$)
@@
OLMo-2-7B-Instruct 0.61 0.38 0.39 0.86 0.84 0.10 0.64 0.61
Table 12: Combined forecasting performance for cutoff 2021-12-01. CRPS (T) denotes the Continuous Ranked Probability Score for Timeframe Prediction, while CRPS (Q) denotes the Continuous Ranked Probability Score for Quantity Estimation.
Model Accuracy ($\uparrow$) F1 ($\uparrow$) Brier ($\downarrow$) ADE ($\downarrow$) CRPS (T) ($\downarrow$) APE ($\downarrow$) MAE ($\downarrow$) CRPS (Q) ($\downarrow$)
@@
OLMo-2-7B-Instruct 0.60 0.33 0.40 0.93 0.90 0.10 0.65 0.62
Table 13: Combined forecasting performance for cutoff 2022-08-01. CRPS (T) denotes the Continuous Ranked Probability Score for Timeframe Prediction, while CRPS (Q) denotes the Continuous Ranked Probability Score for Quantity Estimation.
Model Accuracy ($\uparrow$) F1 ($\uparrow$) Brier ($\downarrow$) ADE ($\downarrow$) CRPS (T) ($\downarrow$) APE ($\downarrow$) MAE ($\downarrow$) CRPS (Q) ($\downarrow$)
@@
OLMo-2-7B-Instruct 0.59 0.33 0.41 0.91 0.87 0.12 0.61 0.57
Table 14: Combined forecasting performance for cutoff 2023-03-01. CRPS (T) denotes the Continuous Ranked Probability Score for Timeframe Prediction, while CRPS (Q) denotes the Continuous Ranked Probability Score for Quantity Estimation.
Model Accuracy ($\uparrow$) F1 ($\uparrow$) Brier ($\downarrow$) ADE ($\downarrow$) CRPS (T) ($\downarrow$) APE ($\downarrow$) MAE ($\downarrow$) CRPS (Q) ($\downarrow$)
GPT2 0.59 0.31 0.37 1.00 1.00 0.25 0.73 0.72
GPT2-XL 0.71 0.30 0.42 1.00 1.00 0.04 0.69 0.64
Pythia-14m 0.62 0.14 0.65 1.00 1.00 0.25 0.85 0.83
Pythia-160m 0.67 0.31 0.51 1.00 1.00 0.02 0.67 0.65
Pythia-2.8b 0.52 0.29 0.45 1.00 1.00 0.12 0.73 0.69
Bloom-560m 0.52 0.38 0.41 1.00 1.00 0.03 0.67 0.65
Bloom-7b1 0.68 0.36 0.30 1.00 1.00 0.11 0.78 0.75
Llama-7b 0.54 0.27 0.55 0.96 0.94 0.06 0.62 0.59
OLMo-1B 0.23 0.21 0.75 1.00 1.00 0.22 0.76 0.74
OLMo-7B 0.22 0.23 0.83 1.00 1.00 0.22 0.82 0.81
OLMo-7B-Instruct 0.68 0.16 0.36 1.00 1.00 0.27 0.81 0.78
OLMo-2-7B 0.51 0.24 0.44 1.00 1.00 0.02 0.65 0.62
OLMo-2-7B-Instruct 0.59 0.30 0.41 0.87 0.82 0.09 0.60 0.57
Table 15: Combined forecasting performance for cutoff 2023-12-01. CRPS (T) denotes the Continuous Ranked Probability Score for Timeframe Prediction, while CRPS (Q) denotes the Continuous Ranked Probability Score for Quantity Estimation.

### v4

@@ 5.3 Results
Table 4 and Table 5 show that while some models reach reasonable accuracy on binary classification tasks, their confidence calibration remains inconsistent. For instance, GPT2-XL achieves a relatively strong accuracy of 0.67 but has a higher Brier score of 0.45, whereas OLMo-7B-Instruct achieves a similar accuracy (0.65) with a lower Brier score of 0.38, indicating better alignment between predicted probabilities and actual outcomes. This discrepancy suggests that models may differ not only in their ability to identify the correct answer, but also in how they distribute confidence over binary options.
Importantly, model performance, whether in terms of accuracy or confidence calibration, does not appear to correlate strongly with scale or training data recency. For example, LLaMA-7B, a more recent model, shows lower accuracy (0.57) and higher Brier score (0.52) than older models like GPT2-XL. Moreover, LLaMA-3.1-8B fails entirely on this task, with an accuracy of 0.08 and a Brier score of 0.99. Manual inspection reveals that it often outputs malformed answers (e.g., long strings of zeros) rather than valid "Yes"/"No" responses, suggesting a failure to follow the prompt format.
@@
Table 4 and Table 5 demonstrate that timeframe prediction is especially difficult for current models. Across the board, both ADE and CRPS values remain close to 1.0, indicating little improvement over naive or uninformed baselines. Only OLMo-2-7B-Instruct achieves a CRPS noticeably below this range (0.82), but even this is far from well-calibrated. The generally poor performance suggests that temporal forecasting presents challenges beyond simple factual retrieval or classification. This difficulty likely reflects the inherent complexity of temporal prediction, which often requires models to reason over event dynamics, causality, and elapsed time, all of which are underrepresented in standard pretraining corpora, as well as the challenge posed by the broader space of possible answers, which extends far beyond binary classifications. In contrast to Boolean tasks, where resolution criteria are clearly defined and outcomes are binary, timeframe questions require mapping language input to a continuous temporal target, which demands a finer-grained understanding of how events evolve over time. That most models fail to depart meaningfully from baseline-level CRPS indicates that LLMs may lack sufficient temporal inductive biases or exposure to structured time-referenced reasoning.
@@
As shown in Table 4 and Table 5, quantity estimation which predicts a numerical outcome exhibits a broader range of performance across models. Some models such as Pythia-160M and GPT2-XL achieve very low absolute percentage error (APE = 0.02 and 0.04 respectively), indicating strong point prediction capability. However, their CRPS values (0.65 and 0.72) suggest relatively poor calibration. In contrast, OLMo-2-7B-Instruct, with a higher APE of 0.08, attains the lowest CRPS (0.57), suggesting it better captures uncertainty even if its point estimates are less precise. This divergence illustrates that point prediction accuracy and confidence calibration are not necessarily aligned. Predicting a correct numerical value does not guarantee that a model’s probability distribution around its prediction reflects uncertainty appropriately. Since CRPS penalizes both distance from the true value and over/under-confidence, this metric provides a more comprehensive view of performance in numerical forecasting tasks. These results suggest that even relatively strong predictive models may require additional mechanisms such as structured uncertainty modeling to produce well-calibrated forecasts in quantity estimation settings.
5.4 Findings
@@
Model size alone does not reliably predict forecasting performance across tasks. As shown in Table 5, within the Pythia family, Pythia-2.8b occasionally improves over smaller variants in certain metrics. For example, it achieves lower CRPS than Pythia-14m on quantity estimation but these gains are neither large nor consistent. In some cases, smaller models like Pythia-160M outperform their larger counterparts in point prediction metrics such as APE. A similar trend appears within the OLMo family: while the OLMo-2-7B variants perform competitively in calibration metrics (e.g., CRPS), this is not accompanied by uniform improvements in accuracy or absolute error. These observations suggest that increases in parameter count do not automatically translate to improved forecasting performance. While larger models often benefit from greater representational capacity, the forecasting task, particularly when it involves expressing well-calibrated uncertainty, may place demands that are orthogonal to model size. Specifically, forecasting tasks may require targeted reasoning over time, quantities, or epistemic uncertainty, which are not clearly emphasized during standard pretraining or scaled parameter growth. These findings indicate that effective forecasting may depend more on model behavior during decoding or task-specific adaptation, rather than scale alone.
Model Acc. ($\uparrow$) F1 ($\uparrow$) Brier ($\downarrow$) ADE ($\downarrow$) CRPS (T) ($\downarrow$) APE ($\downarrow$) MAE ($\downarrow$) CRPS (Q) ($\downarrow$)
@@
Table 6 compares base and instruct-tuned variants of OLMo-7B and OLMo-2-7B across Boolean, timeframe, and quantity forecasting tasks. For Boolean questions, OLMo-7B-Instruct achieves higher accuracy (0.65 vs. 0.21) and a lower Brier score (0.38 vs. 0.82), indicating better confidence calibration. In timeframe prediction, the instruct-tuned OLMo-2-7B-Instruct improves uncertainty estimation, with an ADE of 0.87 and CRPS of 0.82, compared to 0.9981 and 0.9970 (both rounded to 1.00) for the base OLMo-2-7B. For quantity estimation, instruct-tuned models have slightly higher APE but lower MAE and CRPS, suggesting better uncertainty calibration. While these results point to potential benefits of instruction tuning for probabilistic reasoning, the sample size remains limited due to the recency of instruction-tuned models, and the observed gains may not be statistically robust.
Table 7: Different aggregation methods for extracting predictions and confidence from Llama-7B.
Aggregation Method N Acc. ($\uparrow$) F1 ($\uparrow$) Brier ($\downarrow$)
@@
Table 7 compares different aggregation methods for deriving the final prediction and confidence estimate from the top 10 outputs of Llama-7B. Bayesian Aggregation achieves the highest accuracy (0.58) and F1 score (0.35), suggesting it is the most effective at identifying correct outcomes. However, Weighted Average yields a significantly lower Brier score (0.29), indicating superior confidence calibration compared to others. Majority Vote, Highest Confidence, and Logit Mean Probability produce comparable accuracy and F1 scores but have noticeably higher Brier scores, suggesting weaker uncertainty estimation. These results highlight that even when point prediction performance is similar, aggregation methods substantially impact confidence reliability. The challenge remains in developing techniques that optimize both accuracy and calibration simultaneously, emphasizing the importance of uncertainty-aware forecasting. In preliminary experiments, we also explored verbalized confidence (Xiong et al., 2023), prompting the model to explicitly generate a probability (e.g., “I’m 80% confident…”). However, most models either failed to produce enough meaningful numeric estimates or generated inconsistent responses, particularly on harder questions. This highlights a limitation in current instruction-following capabilities when it comes to quantifying uncertainty directly. Therefore, future work includes fine-tuning for calibrated confidence and integrating task-specific priors. Joint modeling of accuracy and uncertainty remains a key challenge.
@@
This observation suggests that forecasting is meaningfully different from retrieval-based tasks such as fact-checking or question answering (Guo et al., 2022). While those tasks often benefit from training data that includes the target facts or related content, forecasting involves reasoning under uncertainty about events that had not yet occurred during training. Models are not explicitly trained to perform such reasoning, and pretraining objectives generally do not reward predictions about future developments. The absence of a correlation between cutoff proximity and performance highlights that forecasting requires capabilities beyond knowledge access, including abstraction, probabilistic reasoning, and temporal inference. These factors make the task more complex and reinforce the need for dedicated methods that go beyond static knowledge retrieval. These findings would not be observable under prior benchmarks that lack confidence metrics or diverse forecasting tasks.
@@ 7 Limitations
FOReCAst is built entirely from Metaculus, a high-quality forecasting platform with clear resolution criteria and an active user base. This provides a strong foundation for defining and benchmarking probabilistic forecasting with LLMs. However, it does not reflect the full diversity of forecasting formats, domains, or styles seen in platforms like Good Judgment Open or domain-specific settings. Expanding to other platforms is an important next step to test generalization across forecasting cultures. Our scoring formulation aligns confidence with meaningful deviations from uncertainty and is tailored to Metaculus conventions. However, it can be adapted for use with other platforms or scoring protocols. The goal is to maintain an interpretable and consistent signal across all tasks.
Gold confidence labels are derived from community-aggregated forecasts, offering a practical proxy for crowd belief and uncertainty. While this enables scalable evaluation, it may obscure forecaster disagreement and introduce biases from group dynamics or correlated errors. Future work could incorporate richer uncertainty signals, such as forecast distributions, forecaster rationales, or disagreement metrics. Finally, FOReCAst focuses on English-language questions. While this is consistent with Metaculus and existing NLP resources, probabilistic reasoning and uncertainty expression can differ across languages and cultures. Extending the benchmark to multilingual settings and modeling temporal adaptation and real-time reasoning are key directions for improving robustness and real-world applicability.
@@ Appendix E Additional Results
Table 10: Combined forecasting performance for cutoff date 2017-12. CRPS (T) denotes the Continuous Ranked Probability Score for Timeframe Prediction, while CRPS (Q) denotes the Continuous Ranked Probability Score for Quantity Estimation.
Model Acc. ($\uparrow$) F1 ($\uparrow$) Brier ($\downarrow$) ADE ($\downarrow$) CRPS (T) ($\downarrow$) APE ($\downarrow$) MAE ($\downarrow$) CRPS (Q) ($\downarrow$)
@@
OLMo-2-7B-Inst 0.59 0.42 0.40 0.84 0.83 0.13 0.69 0.67
Qwen-1.5B-Inst 0.67 0.32 0.32 1.00 1.00 0.07 0.73 0.71
Qwen-7B-Inst 0.68 0.26 0.26 1.00 1.00 0.08 0.63 0.61
Llama-3.1-8B 0.05 0.02 0.96 1.00 1.00 0.03 0.70 0.68
Table 11: Combined forecasting performance for cutoff date 2020-03. CRPS (T) denotes the Continuous Ranked Probability Score for Timeframe Prediction, while CRPS (Q) denotes the Continuous Ranked Probability Score for Quantity Estimation.
Model Acc. ($\uparrow$) F1 ($\uparrow$) Brier ($\downarrow$) ADE ($\downarrow$) CRPS (T) ($\downarrow$) APE ($\downarrow$) MAE ($\downarrow$) CRPS (Q) ($\downarrow$)
@@
OLMo-2-7B-Inst 0.59 0.41 0.40 0.83 0.82 0.12 0.70 0.68
Qwen-1.5B-Inst 0.66 0.32 0.33 1.00 1.00 0.07 0.75 0.73
Qwen-7B-Inst 0.68 0.27 0.26 1.00 1.00 0.07 0.64 0.62
Llama-3.1-8B 0.05 0.02 0.96 1.00 1.00 0.02 0.71 0.69
Table 12: Combined forecasting performance for cutoff date 2021-12. CRPS (T) denotes the Continuous Ranked Probability Score for Timeframe Prediction, while CRPS (Q) denotes the Continuous Ranked Probability Score for Quantity Estimation.
Model Acc. ($\uparrow$) F1 ($\uparrow$) Brier ($\downarrow$) ADE ($\downarrow$) CRPS (T) ($\downarrow$) APE ($\downarrow$) MAE ($\downarrow$) CRPS (Q) ($\downarrow$)
@@
OLMo-2-7B-Inst 0.61 0.38 0.39 0.86 0.84 0.10 0.64 0.61
Qwen-1.5B-Inst 0.66 0.24 0.33 1.00 1.00 0.10 0.75 0.72
Qwen-7B-Inst 0.68 0.14 0.26 1.00 1.00 0.08 0.60 0.59
Llama-3.1-8B 0.06 0.00 0.95 1.00 1.00 0.02 0.62 0.59
Table 13: Combined forecasting performance for cutoff date 2022-08. CRPS (T) denotes the Continuous Ranked Probability Score for Timeframe Prediction, while CRPS (Q) denotes the Continuous Ranked Probability Score for Quantity Estimation.
Model Acc. ($\uparrow$) F1 ($\uparrow$) Brier ($\downarrow$) ADE ($\downarrow$) CRPS (T) ($\downarrow$) APE ($\downarrow$) MAE ($\downarrow$) CRPS (Q) ($\downarrow$)
@@
OLMo-2-7B-Inst 0.60 0.33 0.40 0.93 0.90 0.10 0.65 0.62
Qwen-1.5B-Inst 0.68 0.25 0.31 1.00 1.00 0.11 0.76 0.74
Qwen-7B-Inst 0.68 0.12 0.25 1.00 1.00 0.10 0.60 0.58
Llama-3.1-8B 0.06 0.00 0.96 1.00 1.00 0.02 0.62 0.58
Table 14: Combined forecasting performance for cutoff date 2023-03. CRPS (T) denotes the Continuous Ranked Probability Score for Timeframe Prediction, while CRPS (Q) denotes the Continuous Ranked Probability Score for Quantity Estimation.
Model Acc. ($\uparrow$) F1 ($\uparrow$) Brier ($\downarrow$) ADE ($\downarrow$) CRPS (T) ($\downarrow$) APE ($\downarrow$) MAE ($\downarrow$) CRPS (Q) ($\downarrow$)
@@
OLMo-2-7B-Inst 0.59 0.33 0.41 0.91 0.87 0.12 0.61 0.57
Qwen-1.5B-Inst 0.65 0.20 0.34 1.00 1.00 0.13 0.72 0.69
Qwen-7B-Inst 0.68 0.12 0.25 1.00 1.00 0.10 0.56 0.54
Llama-3.1-8B 0.07 0.00 0.96 1.00 1.00 0.02 0.59 0.55